It is 395 km from ski valley to Vancouver. Sam made the trip in 6h travelling by bus and train. The train averaged 70 km/h and the bus 60 km/h. How much time was spent travelling by the train?
Answer by checkley75(3666)

70X+60(6-X)=395
70X+360-60X=395
10X=395-360
10X=35
X=35/10
X=3.5 HOURS @ 70 KM TRAIN RIDE.
6-3.5=2.5 HOUR @ 60 KM BUS RIDE.
PROOF
70*3.5+60*2.5=395
245+150=395
395=395